Assume an economy produces only footballs and baseballs and the base year is 2016.

Quantities Produced Prices
2016 2017 2016 2017
Hamburgers 200 300 $20 $3.00
Hot dogs 500 600 $10 $1.50

Given the data in the table above, what is the value of real GDP in 2017?

Multiple Choice

  • $9,000

  • $10,000

  • $12,000

  • $12,500
Answer #1

Real GDP in 2017= output of 2017 * price of 2016

Real GDP in 2017=300*20+600*10=6000+6000=12000

ans is C